OpenStudy (anonymous):

what's the equation for a hyperbola with foci at (0, ±10) and vertices at (0, ±8)?

OpenStudy (anonymous):

\[c =10\\ a=8\\ b^2 =c^2 -a^2= 100-64=36\\ \] Can you finish it now?

OpenStudy (anonymous):

\[ \frac {y^2}{64} - \frac {x^2}{36} =1 \]